# Arrows of Time for Large Language Models

### v2

@@ 3.2.3 Experimental Results
In this section, we present experimental results supporting the claims of the previous sections. We first consider a dataset made of linear languages with $x,y\in\mathbb{F}_{2}^{25}$, for different sparsities of $f^{\rightarrow}$. We train a GPT-1 model on these datasets (see Appendix B.2) and plot the final losses in Fig. 3, confirming that sparser matrices are easier to learn.
@@
In the second experiment, we consider a model’s ability to learn a sparse update, given a learned prior: we first train FW/BW models on a linear language with $m=20$ (with a sparse FW matrix) until it is learned perfectly, then generate a new linear language by a sparse perturbation of the learned language. Table 5 shows the losses of both models after 400 gradient descent steps. Again, we see that the FW model adapts better to the sparse modifications (see Appendix B.2.3).
Table 5: GPT-1 losses for different perturbations of the learned prior.
@@ 4 Discussion
In this paper, we have investigated the abilities of auto-regressive LLMs, which predict tokens sequentially: for a given measure (dataset), we compare the abilities of two models (FW/BW). Theoretically, if both models learn to represent the same measure, their average log-perplexities coincide. Somehow surprisingly, we discover the existence of an Arrow of Time (AoT) for natural language datasets: across a wide variety of models and hyperparameters, we see that all else being equal, FW models consistently exhibit a lower perplexity than BW ones; we see this difference emerge as soon as the model is large enough; its causes appear to lie in long-range correlations in the data, as the magnitude of this effect increases with the context length. We then propose a framework to explain this phenomenon, based on computational complexity and sparsity ideas: we construct examples of synthetic datasets based on arithmetic operations that display an asymmetry in terms of computability (despite being information-theoretically reversible); finally, we propose a setup where a FW AoT like the one seen in natural language can spontaneously emerge. Our work suggests a number of possible future research directions:
@@
  • Can AoTs be revealed in continuous settings, e.g. for video prediction?
@@ Acknowledgements
The authors would like to thank Stéphane d’Ascoli, Gloria Capano, Diego Dorn, Franck Gabriel, Ron Maimon, Jacob Menick, Christos Papadimitriou, João Penedones, Matthieu Wyart, Nicolas Zlatoff for interesting discussions. Support from the Blavatnik Family Foundation, the Latsis Foundation, the NCCR SwissMAP, and from an EPFL FSB Seed Funding Grant are gratefully acknowledged.
@@ Appendix A Details on training on natural languages
For any experiment, the precise hyperparameters that were used can be found in the supplementary material, under the folder “Training Parameters”, in .$json$ format. Those files can also be used to reproduce any experiment using the codebase, as explained in the README.md of the Github repo.
All experiments (save for the 512 context size) were run on a single A100 GPU, adjusting the batch size to fit in memory.
Concerning the shuffling of the dataset, we proceed as follows for all the experiments. We begin by separating the textual dataset in ‘sentences’ of the appropriate context size $n$, with a stride of $n/2$ (i.e., if we have a context size of 4 and the text is ABCDEF, this results in two sentences, ABCD and CDEF). This is to ensure that all tokens appear in the training data with at least some context. After that, we shuffle the obtained sentences with a set seed. We withhold 250k sentences (1000 batches at 250 batch size) for validation. The inversion of the tokens is made at the level of each batch; in this way, when training, the FW/BW models see the data in the same order, preventing the emergence of undesirable differences.
@@ A.3 Context Window size
For testing for the influence of the context window length, we use a GPT2-Medium model, with a context window length going from $16$ to $512$ tokens. Because of the different context lengths, it will take a model with a small context length many more gradient steps to see the same amount of data. For this reason, we do not train all models up to the equivalent of 1 epoch of the French dataset ($\sim 26.6B$ tokens), but rather train them for sufficiently long so that perplexity differences stabilize, and that their losses converge. Due to the cosine annealing learning rate schedule, we stop the training at the end of a cosine decay, to avoid the bump in the loss caused by a warm restart (note however that this has little to no effect on the perplexity differences).
@@
Table 7: Number of batches seen during training for the different context lengths. Not that the recorded batch size is “effective”, potentially obtained through aggregation of smaller batch sizes.
@@
For this experiment, because of our memory limitations, we trimmed down the English dataset, which was too big when working with context windows of lengths $16$ and $32$. Note that the 256 context length experiment in this section is thus slightly different from the one recorded in Table 3, due to the different datasets (as well as batch size and cosine annealing period).
@@ A.4 Other Models
The GRU and LSTM implementations we used for these experiments are those natively implemented in the pytorch.nn of the Pytorch Python package. In each case, we choose the “input size” (i.e., the number of embedding dimensions for each token) to be equal to the “hidden size” (i.e., the number of hidden dimensions in the RNN’s hidden state).
@@ A.5 Other Languages
For the experiments on other languages, we decided to stop the training at the equivalent of 1 epoch of Greek, which was one of the smallest datasets on which we were training (along with Turkish). This choice was made so for all languages, models never saw the same datapoint more than once (note that we trained on Greek for two epochs, as shown in Fig. 5; this suggests that our results remain valid beyond one epoch).
For GPT2-Medium, a batch size of $90$ was used, with one warm restart during training. For the XL models, a batch size of $26$ was used, aggregated $6$ times for an effective batch size of $156$, which didn’t allow for a warm restart before the one epoch of Greek.
In Fig. 4 we showcase the validation loss of all the language during training, displaying the same behaviour as observed for English and French in section (2.2.1).
Figure 4: Validation losses through training of GPT2-Medium for different languages, both in FW and BW. Note that we clipped the beginning of training to better showcase the subtle differences in loss.
@@ A.6 BPE Tokenization
For the tokenization, we use the Huggingface implementation of the BPE Tokenizer (link), using the method Tokenizer.train_from_iterator. The tokenizers are trained on the same CC100 dataset on which we train the model.
For the experiment in which we train the BPE tokenizer in reverse (see Section 2.2.6), we first reverse the Greek dataset at the character level (not at the byte level, as this would make the output of the model unreadable), then train the BPE tokenizer on this new dataset. We then train a FW and a BW model on this character-flipped dataset, tokenized with the new BPE tokenizer. We will call a model ‘backward’ (BW) if it processes tokens in the opposite order w.r.t. the natural reading direction (hence the ‘previous-token predictor on the character-flipped dataset’ corresponds to what we will call the FW model). Thus, if the arrow of time is a property of the language and not a tokenization artifact (as we expect), we expect both BW and FW models to perform similarly, independently of the tokenization scheme. Fig. 6 confirms this; relying on the reverse BPE tokenization introduces almost no difference, except for a very small discrepancy for the BW models. Since it produces no noticeable difference for the FW model, this small difference in the BW case is likely due to initialization.
Figure 6: Validation curves for models training on the Greek dataset on forward and backward (label prefixed with ‘rev’) BPE tokenizations. The arrow of time remains in the same, forward direction, despite the different tokenization schemes.
@@ B.1 Matrix Inverse Sparsity
Figure 7: Plot displaying the connection between the sparsity of a matrix and its inverse. It is clear that on average, the inverse of a sparse matrix is less sparse. The number of non-zero elements for the inverse caps at 450450~{}450450
@@ B.2.2 Sparsity Levels
In the first experiment, we generate a linear language model in $\mathbb{F}^{25}_{2}$, with $p=0$ perturbations, and matrices with a number $25+k$ of non-zero elements, where $k\in[0,2,4,8,10,14,18,20,25,30,35,40,45,50]$. We then train a transformer model of size GPT1 (see Table 6) on $600k$ sentences, with batch size $200$. Note that the context size of the model matches exactly the number of tokens in one sentence. Final losses are reported in 3. Note that for lower sparsities, the trend is not obvious: this is due to the high variance in the final learning rate, as the learning of only a few non-zero elements is binary, depending on the initialization, the model either learns the matrix perfectly quickly, or it usually struggles to find the last few non-zero elements. Fig. 8 displays this behavior in the case $k=4$. Note that the perturbations somewhat reduce this behavior, but don’t cancel it completely.
@@ B.2.3 Sparse updates
Figure 10: Training loss for the FW and BW models, trying to learn a linear language with a matrix $k=6$. While the FW model learns the quasi-optimal solution very quickly, the BW model remains stuck on a plateau for a long time. This corresponded to a single element of the predicted vector which was missing, as in Fig. 8.
@@
Figure 11: Averaged loss for forward and backward models, when trying to learn a sparse forward perturbation of the Linear language. In the first $\sim 100$ steps, the curves are similar as they simply reduce their confidence on the new, perturbed tokens to random chance. Subsequently, they begin learning the perturbation. The forward model is clearly at an advantage.

### v3

@@ 3.2.3 Experimental Results
In this section, we present experimental results supporting the claims of the previous sections. We first consider a dataset made of linear languages with $x,y\in\mathbb{F}_{2}^{25}$, for different sparsities of $f^{\rightarrow}$. We train a GPT1 model on these datasets (see Appendix B.2) and plot the final losses in Fig. 4, confirming that sparser matrices are easier to learn.
@@
In the second experiment, we consider a model’s ability to learn a sparse update, given a learned prior: we first train FW/BW models on a linear language with a $m=20$ sparse FW matrix, until it is learned perfectly, then generate a new linear language by a sparse perturbation of the matrix. Table 5 shows the losses of both models after 400 gradient descent steps. Again, we see that the FW model adapts better to the sparse modifications (see also Appendix B.2.3).
Table 5: GPT losses for various perturbations of the learned prior.
@@ 4 Discussion
In this paper, we have investigated the abilities of auto-regressive LLMs, which predict tokens sequentially: for a given measure (dataset), we compare the abilities of two models (FW/BW). Theoretically, if both models learn to represent the same measure, their average log-perplexities should coincide. We discover the existence of an Arrow of Time (AoT) for natural language datasets: across a wide variety of models and hyperparameters, all else being equal, FW models exhibit a consistently lower perplexity than BW ones; this difference emerges as soon as the model is large enough; its causes appear rooted in long-range correlations in the data, as the effect magnitude increases with the context length. We propose a framework to explain this phenomenon, based on complexity and sparsity ideas: we construct examples of synthetic datasets based on operations that display an asymmetry in terms of computability (despite being information-theoretically reversible); finally, we propose a setup where a FW AoT like the one seen in natural language can spontaneously emerge. Our work suggests a number of possible future research directions:
@@
  • Can AoTs, Causality, and $\mathcal{V}$-Information (see Section 1.6) be understood under a common framework?
  • What about AoTs in continuous settings, e.g. for video?
@@
  • Is the presence of an AoT in data a sign of life or intelligent processing?
@@ Acknowledgements
The authors would like to thank Stéphane d’Ascoli, Samy Bengio, Gloria Capano, Diego Dorn, Franck Gabriel, Ron Maimon, Jacob Menick, Christos Papadimitriou, João Penedones, Matthieu Wyart, Nicolas Zlatoff, as well as the anonymous reviewers for interesting discussions and suggestions. Support from the Blavatnik Family Foundation, the Latsis Foundation, the NCCR SwissMAP, and from an EPFL FSB Seed Funding Grant are gratefully acknowledged.
Impact Statement
This paper presents work with the goal of advancing the field of Machine Learning, and our scientific understanding of language models. There are many potential societal consequences of a greater understanding of this discipline, and thus, indirectly, of our work, however, none of them feel direct enough to be specifically highlighted here.
@@ Appendix A Details on training on natural languages
For any experiment, the precise hyperparameters used can be found in the code repository found at github.com/frotaur/ICMLBackPerp, under the folder ‘Training Parameters’, in .$json$ format. Those files can also be used to reproduce any experiment using the codebase, as explained in the README.md of the repository The branch ‘main’ contains the code used to generate the experiments of the paper, while the ‘Rebutt’ branch contains updated, cleaner code, as well as the necessary data to run the extra experiments on Tagalog, Hebrew and Arabic.
All experiments (save for the 512 context size) were run on a single A100 GPU, adjusting the batch size to fit the available memory.
Concerning the shuffling of the dataset, we proceed for all the experiments as follows: we begin by splitting the textual dataset into ‘sentences’ of the appropriate context size $n$, with a stride of $n/2$ (i.e., if we have a context size of 4 and the text is ABCDEF, this results in two sentences, ABCD and CDEF). This is to ensure that all tokens appear in the training data with at least some context. After that, we shuffle the obtained sentences with a set seed. We withhold 250k sentences (1000 batches at batch size 250) for validation. The inversion of the tokens is made at the level of each batch; in this way, when training, the FW/BW models see the data in the same order, preventing the emergence of undesirable differences.
@@ A.3 Context Window size
For the testing of the influence of the context window length, we use a GPT2-Medium model, with context window lengths going from $16$ to $512$ tokens. Because of the different context lengths, it will take a model with a small context length many more gradient steps to see the same amount of data. For this reason, we do not train all models up to the equivalent of 1 epoch of the French dataset ($\sim 26.6B$ tokens), but rather train them for sufficiently long so that the perplexity differences stabilize, and that their losses converge. Due to the cosine annealing learning rate schedule, we stop the training at the end of a cosine decay, to avoid the bump in the loss caused by a warm restart (note however that this has little to no effect on the perplexity differences).
@@
Table 7: Number of batches seen during training for the different context lengths. Note that the recorded batch size is the ‘effective’ one, that is, potentially obtained through aggregation of smaller batch sizes.
@@
For this experiment, because of our memory limitations, we trimmed down the English dataset, which was too big when working with context windows of lengths $16$ and $32$. Note that the 256 context length experiment in this section is thus slightly different from the one recorded in Table 3, due to the different datasets (as well as the batch size and the cosine annealing period).
@@ A.4 Other Models
The GRU and LSTM implementations we used for these experiments are those natively implemented in the pytorch.nn module of the Pytorch Python library. In each case, we choose the ‘input size’ (i.e., the number of embedding dimensions for each token) to be equal to the ‘hidden size’ (i.e., the number of hidden dimensions in the RNN’s hidden state).
@@ A.5 Other Languages
For the experiments on other languages, we decided to stop the training at the equivalent of 1 epoch of Greek, which was one of the smallest datasets on which we were training (along with Turkish). This choice was maintained for all languages, and models never saw the same datapoint more than once (note that we also tried training on Greek for two epochs, as shown in Fig. 5; this suggests that our results remain valid beyond one epoch).
For GPT2-Medium, a batch size of $90$ was used [^1] Due to an oversight in the code, the batches were not aggregated in groups of 2 as expected, but the loss was still renormalized by dividing it by 2. This amounts to a very slight change in learning dynamics, but does not affect any of the results. Similarly, all graphs/reported results display the correct loss. In case one wants to reproduce exactly the results of the paper, the loss should be divided by 2 before backpropagation., with one warm restart during training. For the XL models, a batch size of $26$ was used, aggregated $6$ times for an effective batch size of $156$, which didn’t allow for a warm restart before the one epoch of Greek.
We also tested (using GPT2-Medium, with an adjusted learning rate schedule, due to the smaller size of the datasets) Hebrew, Arabic and Tagalog (see Table 9 for final losses), at the suggestions of anonymous reviewers.
Hebrew and Arabic constituted an example of languages written right-to-left; a priori, we would not expect this to affect the emergence of an AoT: after all, tokens are still processed in the ‘spoken’ order by the model, so the writing direction does not affect training. Still, there could have been an influence on the language itself from the writing direction, which cannot be detected from our results.
Tagalog is an example of a language with ‘verb-initial word order’, a relatively rare class of languages, which was not included in our list. Here, our expectations are again confirmed as an AoT appears also for this example. This reinforces the idea that the AoT for natural language emerges from long-range correlations. The specifics of the grammar and the order of words in a sentence are therefore not that important.
Table 9: Final losses for extra languages. Arabic and Hebrew are at 1 epoch of training, and Tagalog at 7 epochs, due to the small size of the dataset. Format: [Final FW loss]/[BW relative difference].
Tagalog Arabic Hebrew
Med 2.368/+1.48% 3.446/+1.91% 3.288/+2.37%
@@ A.6 BPE Tokenization
For the tokenization, we use the Huggingface implementation of the BPE Tokenizer (link), using the method Tokenizer.train_from_iterator. The tokenizers are trained on the same CC-100 dataset on which we train the model.
To exclude potential tokenization asymmetries (see section 2.2.6), we perform extra experiments in which we train the BPE tokenizer in reverse. To do so, we reverse the language dataset at the character level (not at the byte level, as this would make the output of the model unreadable because of multi-byte characters in utf-8), then train the BPE tokenizer on this new dataset. We then train a FW and a BW model on this character-flipped dataset, tokenized with the new BPE tokenizer.
To make things clearer, we will call a model ‘backward’ (BW) if it processes tokens in the opposite order w.r.t. the natural reading direction (hence the ‘previous-token predictor’ on the ‘character-flipped dataset’ corresponds to what we will call the FW model). Thus, if the Arrow of Time is a property of the language and not a tokenization artifact (as we expect), we expect the arrow of time to be in the same direction, independently of the tokenization scheme. Fig. 6 confirms this; relying on the reverse BPE tokenization introduces very slight differences in the losses. This difference is negligible compared to the AoT effect in both Greek and French, so we can conclude that the AoT is not a tokenization artifact.
In figures 6 and 7, we display the loss curves during training for the french and greek models, trained using both the normal and reversed BPE tokenizers. The tokenizer switch has minimal impact on the losses, and most importantly, it does not affect the AoT direction.
Figure 6: Validation curves for models training on the Greek dataset on forward and backward (label prefixed with ‘rev’) BPE tokenizations. The Arrow of Time remains the same in the FW direction, despite the different tokenization schemes.
Figure 7: Validation curves for models training on the french dataset on forward and backward (label prefixed with ‘rev’) BPE tokenizations. The Arrow of Time remains the same in the FW direction, despite the different tokenization schemes. It seems the reverse tokenization is slightly suboptimal, slightly degrading the losses of both FW and BW models.
A.7 Initialization variance
The AoT computed in our experiments were obtained with a single training run. One might ask if such a difference remains significant compared to variations in loss due to initialization. The consistency throughout experiments shows that the AoT is significant, but in this section we set to verify the magnitude of the initialization variance. Due to computational costs, it is not possible to obtain error bars for all the experiments. Instead, we focus on Greek, for which we re-run the GPT2-Medium training 4 times in total. Computing the error bars, we obtain a loss (at one epoch) of $2.802\pm 0.005$ FW, and $2.871\pm 0.003$ BW. This gives us an AoT magnitude of $2.46\%\pm 0.18\%$, where we can see that the variations due to initialization are negligible w.r.t. the magnitude of the AoT.
@@ B.1 Matrix Inverse Sparsity
Figure 8: Plot displaying the connection between the sparsity of a matrix and its inverse. It is clear that on average, the inverse of a sparse matrix is less sparse. The number of non-zero elements for the inverse caps at 450450~{}450450.
@@ B.2.2 Sparsity Levels
In the first experiment, we generate a linear language model in $\mathbb{F}^{25}_{2}$, with $p=0$ perturbations, and matrices with a number $25+k$ of non-zero elements, where $k\in[0,2,4,8,10,14,18,20,25,30,35,40,45,50]$. We then train a transformer model of size GPT1 (see Table 6) on $600k$ sentences, with batch size $200$. Note that the context size of the model matches exactly the number of tokens in one sentence. Final losses are reported in 4. Note that for lower sparsities, the trend is not obvious: this is due to the high variance in the final learning rate, as the learning of only a few non-zero elements is binary, depending on the initialization, the model either learns the matrix perfectly quickly, or it usually struggles to find the last few non-zero elements. Fig. 9 displays this behavior in the case $k=4$. Note that the perturbations somewhat reduce this behavior, but don’t cancel it completely.
@@ B.2.3 Sparse updates
Figure 11: Training loss for the FW and BW models, trying to learn a linear language with a matrix $k=6$. While the FW model learns the quasi-optimal solution very quickly, the BW model remains stuck on a plateau for a long time. This in fact corresponds to a single element of the predicted vector which was missing, as in Fig. 9.
@@
Figure 12: Averaged loss for forward and backward models, when trying to learn a sparse forward perturbation of the Linear language. In the first $\sim 100$ steps, the curves are similar as both models decrease their confidence in the new, perturbed tokens, setting them back to random chance. Subsequently, they begin learning the perturbation, where the forward model is clearly at an advantage.